Kevin earned $165 interest for 1 year on an investment of $1500. At the same rate, what amount of interest would be earned by an investment of $2500?
I believe the interst rate is 11% which leaves an equation of:
11% * $2500 = $275.00
Is this correct?

Answer by Nate(3500)

Firstly, what is the rate?
Interest = Time * Principal * Rate
165 = 1 * 1500 * r
165 / 1500 = r
At the same rate, what amount of interest would be earned by an investment of $2500?
1*(165/1500)*2500 = Interest after 1 year